OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x2 − 6x − 36

OpenStudy (anonymous):

2x^2 - 6x - 36 -- notice all numbers are even, so you can take 2 out like this: 2(x^2 - 3x - 18) -- now factor... 2(x - 6) (x + 3)

OpenStudy (anonymous):

Which equation is a quadratic equation? 3x − y = 8x + 2 y = 7x(5 − 2x) + 9 7x2 + y = (3 + x)(7x − 1) y − 3 = (x2 − 3)(8x + 3)